OpenStudy (anonymous):

Solve for x. Round your answer to two decimal places. Show your work for full credit. A right triangle is shown with the hypotenuse labeled x. One angle has a measure of thirty nine degrees, and the opposite leg has a length of nine.

OpenStudy (anonymous):

 SOLVE for x

OpenStudy (raden):

use the definiton of : sin(39) = 9/x solve for x and the calculator

OpenStudy (raden):

* definition

OpenStudy (anonymous):

I did that, and my teacher marked me off, and said I have incorrectly solved for x. is there another alternate way?

OpenStudy (raden):

i think only by using trigono be solved

OpenStudy (anonymous):

thank you for your help I will try that

OpenStudy (campbell_st):

ok... so this is a trigonometry question where \[\sin(39) = \frac{9}{x}\] so \[x = \frac{9}{\sin(39)}\] perhaps you needed the step above beofre calculating

OpenStudy (anonymous):

so it would be 9/sin39 not sin(39)/9....?

OpenStudy (campbell_st):

thats correct... sin = opposite/hypotenuse to show you \[\sin(39) = \frac{9}{x}\] multiply both sides by x x sin(39) = 9 divide both sides by sin(39) \[x = \frac{9}{\sin(39)}\]

OpenStudy (anonymous):

That makes much more sense.....I see what you did! but I keep getting a strange decimal....

OpenStudy (campbell_st):

well you are finding the hypotenuse, so the answer needs to larger than the given side of 9. what answer do you have..?

OpenStudy (anonymous):

I got 0.629 which I know is wrong

OpenStudy (campbell_st):

ok... so have a look at you calculator screen... has got a D in the top of the display.... meaning you're working in degrees...

OpenStudy (anonymous):

I am in degrees right now.

OpenStudy (campbell_st):

so you have entered \[9 \div \sin(39) =\]

OpenStudy (anonymous):

yes and it gave me a run on decimal.

OpenStudy (anonymous):

wait wait... i just did it and it gave me 14.30114156 I think that is it!!!

OpenStudy (campbell_st):

thats it

OpenStudy (anonymous):

YAAAAY! thank you so much! I am so grateful! you are awesome